1. Determine the present value of each of three annual coupon payments for a bond with a maturity of four years that pays $500 coupons each year beginning one year from now and has a face value of $20,000. (Assume a current market interest rate of 6%.) What is the current market value of this bond?
Answer #1

Given that: st pv. of nd 2 pr of Coupon = 500/1.06 =471.70 coupon = 500/1.06 = 445.0 coupon = 500/1.063-419.81 pr of zod mar